OpenStudy (anonymous):

how do i do this!? please help! Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2). Answer y = 5/2 x - 2 y = -5/2 x + 8 Y = 2/5 x - 6 y = -2/5 x + 4

OpenStudy (hba):

Well the formula is, \[(y-y_1)=\frac{ y_2-y_1 }{ x_2-x_1 }(x-x_1)\]

OpenStudy (hba):

@Josh45680 Giving direct answers is not allowed on OS. Please check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

anyways.. thank you both.

OpenStudy (hba):

Welcome :)

OpenStudy (anonymous):

To find the answer you should use y = mx + b To find M use the formula: \[\frac{ y2-y1 }{ x2-x1 }\] Then, to find B, use \[b = y1 - mx1\] Finally, put it all together.